### v2

@@ 5.3 Noise affected by a nonlinear combinations of all 100 variables
where for all $y\in\mathbb{Y}$$\pi_{\ell}(y)\triangleq\prod_{i=5\ell+1}^{5(\ell+1)}y_{i}$ for each $\ell\in\llbracket 0,19\rrbracket$ and $f(y)\triangleq\frac{1}{5}\sum_{\ell=0}^{19}\pi_{\ell}(y)$ have explicit definition, and where the blackbox $\varepsilon$ satisfies $\varepsilon(x)\triangleq\exp(\frac{1}{z-4})+\frac{1}{5}\sqrt{\lvert{z-4}\rvert}$ if $z\neq 4$ and $\varepsilon(4)\triangleq+\infty$. Here $\varphi$ has no global minimizer, but $\inf\varphi(\Omega)=0$ and $y^{*}\triangleq\mathds{1}$ is a generalized global solution (even though $\varphi(y^{*})=+\infty$). For all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, we define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$. Then, $\varphi_{|\mathbb{Y}(x)}$ is a smooth function since $\varepsilon(f(y))$ is constant. Moreover its minimizer is analytically obtainable as follows. For all $x\in\mathbb{X}$$\varphi_{|\mathbb{Y}(x)}(y)\geq 0$ and equals 00 when all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ are equal. Under the constraint $y\in\Omega$, equalling all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ implies that $y=\alpha\mathds{1}$ for some $\alpha\in\mathbb{R}$. Then, $y\in\mathbb{Y}(x)$ leads to $4\alpha^{5}=x$. The value of $\alpha$ follows and we get that
@@
Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly. Even with $\tau=100$, the POf provides a gain of performance. NOMAD and PRIMA solving the original problem both fail to significantly improve their initial incumbent solution and, in the second case, both solvers behave as if their initial incumbent solution cannot be improved. Within a comparable budget, our cDSM converges towards the global minimizer on all but one case. Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{ref}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large. Our instance lacks one since it lacks a search step.
Figure 13: Comparison of the best solution found by each solver depending on the computational cost spent. The starting points are chosen as $y^{0}_{1}\triangleq(i/100)_{i=1}^{100}$$y^{0}_{2}\triangleq 0.5\mathds{1}$$y^{0}_{3}\triangleq(\frac{2i}{100})_{i=1}^{100}$, and $y^{0}_{i}$, for each $i\in\llbracket 4,6\rrbracket$, is chosen randomly but such that for each $j\in\llbracket 1,100\rrbracket$, the $j^{th}$ component of $y^{0}_{i}$ lies in $[\frac{j-1}{100},\frac{j}{100}]$.
@@ 5.4 Ten-dimensional noise with non-analytical oracle and 100 variables
$f:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{X}\\ (y_{i})_{i=1}^{100}&\mapsto&\left(g_{1}(y_{10})-\sum\limits_{i=1}^{9}y_{i},\leavevmode\nobreak\ \dots,\leavevmode\nobreak\ g_{10}(y_{100})-\sum\limits_{i=91}^{99}y_{i}\right),\end{array}\right.$
@@
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varepsilon(x)+\sum\limits_{j=1}^{10}\left(\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}\left\lvert{y_{i}}\right\rvert+\left\lvert{g_{j}^{-1}\left(x_{j}+\sum\limits_{i=10(j-1)+1}^{10(j-1)+9}y_{i}\right)}\right\rvert\right).$
@@
$\gamma(x)\triangleq(\gamma_{i}(x))_{i=1}^{100}\quad\mbox{where}\quad\gamma_{i}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ g_{j}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
@@
$\Phi:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{g_{j}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
@@
Unfortunately, $\gamma$ admits no analytical expression since, for all $j\in\llbracket 1,10\rrbracket$$g_{j}^{-1}$ cannot be expressed with elementary functions. However, since $g_{j}^{-1}(x_{j})$ solves the equation $(z+(1+\frac{j}{10})^{z}-1)-x_{j}=0$ (with variable $z\in\mathbb{R}$), we approximately solve this equation using a dichotomic search to obtain an approximate solution $\hat{g_{j}}^{-1}(x_{j})\approx g_{j}^{-1}(x_{j})$. As a result, for all $x\in X$, we approximate
$\gamma(x)=(\gamma_{i}(x))_{i=1}^{100}\approx\hat{\gamma}(x)\triangleq(\hat{\gamma_{i}}(x))_{i=1}^{100}\quad\mbox{where}\quad\hat{\gamma_{i}}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\leavevmode\nobreak\ i\notin 10\mathbb{N},\\ \hat{g_{j}}^{-1}(x_{j})&\mbox{if}\leavevmode\nobreak\ i=10j,\leavevmode\nobreak\ j\in\llbracket 1,10\rrbracket,\end{array}\right.$
@@
$\Phi\approx\hat{\Phi}\quad\mbox{where}\quad\hat{\Phi}:\left\{\begin{array}[]{ccl}\mathbb{X}&\to&\mathbb{R}\\ x\triangleq(x_{j})_{j=1}^{10}&\mapsto&\varepsilon(x)+\sum\limits_{j=1}^{10}\left\lvert{\hat{g_{j}}^{-1}(x_{j})}\right\rvert.\end{array}\right.$
Figure 14 shows a comparison between the different strategies, where for tractability our naive cDSM actually minimizes $\hat{\Phi}$ instead of $\Phi$. Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.15.2 and 5.3. Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of the original problem accordingly. Using our cDSM, the POf provides a gain of performance if we allow $\tau\approx 10$ at most. With $\tau=10$, the POf allows our cDSM to converge under a computational budget similar to whose required by NOMAD and PRIMA, but the solutions returned by our cDSM are $21E$ 2 times better than those returned by NOMAD and $71E$ 7 times better than those returned by PRIMA. The value $\tau=10$ is quite low, but we could consider a higher value by using a better implementation of our algorithm. For example, a test with NOMAD solving the reformulated problem (instead of our naive instance of cDSM) provided similar graphs with $\tau=30$. This last run is represented on Figure 15.
Figure 14: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-5,5]^{100}$.
@@ 6 General discussion
We conclude this work with a discussion on the POf. Section 6.1 summarizes the most important aspects of the POf, and Section 6.2 lists some routes for improvements.
6.1 Comments on the POf
@@
The most stringent aspect of the POf is the oracle function $\gamma$. For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) must have a computable global solution. Ensuring this requirement is challenging in general and likely requires a problem-dependent strategy. Nevertheless, Sections 4.4 and 5.4 show that this theoretical requirement may be slightly relaxed in practice. The POf does not need that $\gamma$ has a closed-form expression, as $\gamma$ may be defined pointwise as, for all $x\in X$, the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with great confidence to approximate closely a global solution. This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.24 and 5), and we are working on possible ways to relax this requirement (see Section 6.2).
Our assumptions are another nontrivial aspect of the POf. Requirements about $\chi$ are presumably easy to satisfy since $\chi$ follows directly from the partition (which is chosen by the user), but we are not aware of any easy and exhaustive rule that ensures Assumption 1 and 2. Nevertheless, they become tractable in some cases. For example, in the context studied in Section 4, Assumptions 1 and 2 hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [2]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets.
A careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous. Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). The POf sometimes leads to an instance where this is possible. A nontrivial example from optimal control theory is given in [9, Chapter 7.3.1].
Theorem 2 is stronger if we moreover assume that $\varphi$ is continuous and $\Omega$ is closed. In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) for all $x^{*}$ (since $\varphi$ is therefore lower semicontinuous and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ naturally holds). In our analysis where this additional assumption does not holds, Theorem 2 states only a sufficient but not necessary condition to claim that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$). Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not lower-semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed by Theorem 2.
@@
Theorem 2 is stated for the cDSM only, but it is actually compatible with many others DFO algorithms. As discussed in Section 3.1, Theorem 2 remains valid if the cDSM in Algorithm 1 is replaced by any DFO algorithm enhanced with a covering step. In general, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is akin to a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods. In particular, if $\dim(\mathbb{X})$ is too large, the cDSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) while another DFO may be more appropriate. We refer to [2] for a guideline showing how to indeed fit a covering into any existing DFO algorithm. However, to our best knowledge, no popular DFO solver implements some algorithms with a covering step yet. Then, in practical settings, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality). Thus Theorem 2 is not yet compatible with most DFO solvers.
@@
The usual case of application of the POf is when $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$. It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) with the cDSM [5, Section 1.4]. Although our examples in Section 5 illustrate cases with $\dim(\mathbb{Y})=101$ at most, we stress that the performance of the POf depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$. Hence $\dim(\mathbb{Y})$ may be possibly infinite, as in [10]. Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be. As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly, then the POf would be deemed useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$.
In Section 4, we show that the POf tackles a class of greybox composite problem. Although the literature on composite problem is rich, [21, 23] claim that most papers assume specific structures for $\widetilde{\varphi}$ and $\sigma$ and $\varepsilon$. We observe that our approach has complementarity with those in [21, 23]. Indeed, [21] considers cases where $\widetilde{\varphi}$ is cheap to evaluate while $\sigma$ and $\varepsilon$ are possibly expensive, but they consider $\widetilde{\varphi}$ with a complex structure and $\dim(\mathbb{Y})$ small and $\dim(\mathbb{X})\leq 65$, while we better tackle cases with $\widetilde{\varphi}$ with simple structure and $\dim(\mathbb{Y})$ large but $\dim(\mathbb{X})$ small. Also, [23] also considers $\widetilde{\varphi}$ well structured and cheap to evaluate, but unlike us it assumes that $\sigma$ and $\varepsilon$ are smooth (although expensive to calculate).

### v3

@@ 5.3 Noise affected by a nonlinear combination of all 100 variables
where for all $y\in\mathbb{Y}$, $\pi_{\ell}(y)\triangleq\prod_{i=5\ell+1}^{5(\ell+1)}y_{i}$ for each $\ell\in\llbracket 0,19\rrbracket$ and $f(y)\triangleq\frac{1}{5}\sum_{\ell=0}^{19}\pi_{\ell}(y)$ have explicit definition, and where the blackbox $\varepsilon$ satisfies $\varepsilon(x)\triangleq\exp(\frac{1}{z-4})+\frac{1}{5}\sqrt{\lvert{z-4}\rvert}$ if $z\neq 4$ and $\varepsilon(4)\triangleq+\infty$. Here $\varphi$ has no global minimizer, but $\inf\varphi(\Omega)=0$ and $y^{*}\triangleq\mathds{1}$ is a generalized global solution (even though $\varphi(y^{*})=+\infty$). For all $x\in\mathbb{X}=X\triangleq\mathbb{R}$, we define $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$. Then, $\varphi_{|\mathbb{Y}(x)}$ is a smooth function since $\varepsilon(f(y))$ is constant. Moreover, its minimizer is analytically obtainable as follows. For all $x\in\mathbb{X}$, $\varphi_{|\mathbb{Y}(x)}(y)\geq 0$ and equals 0 when all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ are equal. Under the constraint $y\in\Omega$, equalling all $(\pi_{\ell}(y))_{\ell\in\llbracket 0,19\rrbracket}$ implies that $y=\alpha\mathds{1}$ for some $\alpha\in\mathbb{R}$. Then, $y\in\mathbb{Y}(x)$ leads to $4\alpha^{5}=x$. The value of $\alpha$ follow, and we get that
@@
Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly. Even with $\tau=100$, the POf provides a gain of performance. NOMAD and PRIMA solving the original problem both fail to significantly improve their initial incumbent solution and, in the second case, both solvers behave as if their initial incumbent solution cannot be improved. Within a comparable budget, the cDSM in our DFPOm converges towards the global minimizer on all but one case. Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{r}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large. Our instance lacks one since it lacks a search step.
Figure 13: Comparison of the best solution found by each solver depending on the computational cost spent. The starting points are chosen as $y^{0}_{1}\triangleq(i/100)_{i=1}^{100}$, $y^{0}_{2}\triangleq 0.5\mathds{1}$, $y^{0}_{3}\triangleq(\frac{2i}{100})_{i=1}^{100}$, and $y^{0}_{i}$, for each $i\in\llbracket 4,6\rrbracket$, is chosen randomly but such that for each $j\in\llbracket 1,100\rrbracket$, the $j^{th}$ component of $y^{0}_{i}$ lies in $\frac{j-1}{100},\frac{j}{100}]$.
@@ 5.4 Ten-dimensional noise with non-analytical oracle and 100 variables
$f:\left\{\begin{array}[]{ccl}\mathbb{Y}&\to&\mathbb{X}\\ (y_{i})_{i=1}^{100}&\mapsto&\left(g_{1}(y_{10})-\sum\limits_{i=1}^{9}y_{i},\penalty 10000\ \dots,\penalty 10000\ g_{10}(y_{100})-\sum\limits_{i=91}^{99}y_{i}\right),\end{array}\right$.
@@
[…]
@@
$\gamma(x)\triangleq(\gamma_{i}(x))_{i=1}^{100}\quad\mbox{where}\quad\gamma_{i}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\penalty 10000\ i\notin 10\mathbb{N},\\ g_{j}^{-1}(x_{j})&\mbox{if}\penalty 10000\ i=10j,\penalty 10000\ j\in\llbracket 1,10\rrbracket,\end{array}\right$.
@@
[…]
@@
Unfortunately, $\gamma$ admits no analytical expression since, for all $j\in\llbracket 1,10\rrbracket$, $g_{j}^{-1}$ cannot be expressed with elementary functions. However, since $g_{j}^{-1}(x_{j})$ solves the equation $(z+(1+\frac{j}{10})^{z}-1)-x_{j}=0$ (with variable $z\in\mathbb{R}$), we approximately solve this equation using a dichotomic search to obtain an approximate solution $\hat{g_{j}}^{-1}(x_{j})\approx g_{j}^{-1}(x_{j})$. As a result, for all $x\in X$, we approximate
$\gamma(x)=(\gamma_{i}(x))_{i=1}^{100}\approx\hat{\gamma}(x)\triangleq(\hat{\gamma_{i}}(x))_{i=1}^{100}\quad\mbox{where}\quad\hat{\gamma_{i}}(x)\triangleq\left\{\begin{array}[]{ll}0&\mbox{if}\penalty 10000\ i\notin 10\mathbb{N},\\ \hat{g_{j}}^{-1}(x_{j})&\mbox{if}\penalty 10000\ i=10j,\penalty 10000\ j\in\llbracket 1,10\rrbracket,\end{array}\right$.
@@
[…]
Figure 14 shows a comparison between the different strategies, where for tractability we actually minimize $\hat{\Phi}$ instead of $\Phi$. Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.1, 5.2 and 5.3. Reformulation ($\mathbf{P}_{\texttt{r}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of Problem ($\mathbf{P}_{\texttt{i}}$) accordingly. Using our cDSM, the DFPOm returns solutions $1E2$ times better than those returned by NOMAD and $1E7$ times better than those returned by PRIMA, and the computational budget is similar if we allow $\tau\approx 10$ at most. The value $\tau=10$ is low, but we could consider a higher value by using a better algorithm. For example, a test with NOMAD solving the reformulated problem (instead of our naive instance of cDSM) provides similar graphs with $\tau=30$, represented on Figure 15.
Figure 14: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$, $i\in\llbracket 1,6\rrbracket$, are chosen as 66 observations of the random uniform independent distribution over $5,5]^{100}$.
@@ 6 General discussion
We conclude this work with a discussion on our contribution. Section 6.1 summarizes the most important aspects of the POf, and Section 6.2 lists some routes for improvements.
6.1 Comments on the POf and the DFPOm
The POf is a reformulation framework that applies to a large class of problems, including nonconvex, discontinuous and unbounded problems. The idea is to pinpoint the difficult aspect in Problem ($\mathbf{P}_{\texttt{i}}$), and to partition the variables space so that this difficulty is transparent in each partition set. Then, the POf defines the oracle function $\gamma$ (assumed computable) that is assumed to return a solution to the problem restricted to any partition set (which is easier to solve than the unrestricted problem). Thus, the POf is not a DFO framework by itself, but rather a way to reformulate a problem with localized difficulties so that Reformulation ($\mathbf{P}_{\texttt{r}}$) tackles only these difficulties. Then, Reformulation ($\mathbf{P}_{\texttt{r}}$) is manageable using DFO algorithms, as we formalize as the DFPOm. This method is theoretically grounded (as formalized by Theorem 2), since Theorem 1 ensures that solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) lead to (possibly, generalized) local solutions to Problem ($\mathbf{P}_{\texttt{i}}$), and Proposition 1 ensures that Algorithm 1 indeed solves Reformulation ($\mathbf{P}_{\texttt{r}}$).
@@ 6.1 Comments on the POf and the DFPOm
A stringent practical aspect of the POf is the oracle function $\gamma$. For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) must have a computable global solution. Ensuring this requirement is challenging and likely problem-dependent. In some cases, $\gamma$ is analytically tractable. However, the POf does not need that $\gamma$ has a closed-form expression. In some of our examples, $\gamma$ is defined pointwise as, for all $x\in X$, the output of some numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) with high confidence to identify a global solution (e.g., by solving a simple smooth subproblem). Our experiments in our former work [11] even show that the DFPOm remains effective when $\gamma(x)$ consists in the output of some solver tackling Subproblem ($\mathbf{P}_{\texttt{s}}(x)$), even if the solver possesses only guarantees to identify a point satisfying necessary optimality conditions. These numerical studies indicate that such approximations are sufficient in practice to reap the benefits of the POf, while our theory justifies the methodology when $\gamma$ returns global solutions. This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.2, 4 and 5), and we are working on possible ways to significantly relax this requirement (see Section 6.2).
@@
Our DFPOm solves Reformulation ($\mathbf{P}_{\texttt{r}}$) using a DFO algorithm with a covering step. We consider the cDSM (Algorithm 2) in our numerical experiments, but any DFO algorithm may be enhanced with a covering step, as discussed in Section 3.2 and [3]. The choice of the most appropriate DFO algorithm is likely to depend on $\Phi$. In general, Reformulation ($\mathbf{P}_{\texttt{r}}$) has little specificity that differentiates it from a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods. In particular, if $\dim(\mathbb{X})$ is too large, a DSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) while another DFO may be more appropriate. However, to our best knowledge, no popular DFO solver implements some algorithms with a covering step yet. Then, in practical settings, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality). Thus, Theorem 2 is not yet compatible with most DFO solvers. Nevertheless, [3] remarks that the covering step is likely optional in practice. We second this observation in Section 5.4, as solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using NOMAD instead of the cDSM provides better results, even though NOMAD does not implement a covering step.
@@
The usual case of application of the DFPOm is $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$. It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{r}}$) with the cDSM [6, Section 1.4]. Although our examples in Section 5 illustrate cases with $\dim(\mathbb{Y})=101$ at most, we stress that the performance of the POf depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$. Hence, $\dim(\mathbb{Y})$ may be possibly infinite, as in [11]. Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be. As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{r}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{i}}$) directly, then the DFPOm is useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$. We also remark that the way the partition is indexed has a direct impact on the behaviour of $\Phi$. The solving process may be substantially simplified by a re-parametrization of $\mathbb{X}$ that avoid concentrations of discontinuities or mitigate local variations of $\Phi$.
For theoretical considerations, our main contributions are Theorems 1 and 2, and Assumptions 1 and 2 which ensure their validity. We did not find generic rules that ensure Assumptions 1 and 2, but nevertheless they are not very demanding in practice, as we discuss in Remarks 2 and 3. Moreover, they become tractable in some cases such as in Section 4. Indeed, they hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [3]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets. Moreover, a careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous. Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{r}}$). This scenario arises in some nontrivial cases, for example from optimal control theory as in [10, Chapter 7.3.1]. We also remark that Theorem 2 is stronger if we moreover assume that $\varphi$ is lower semicontinuous and $\Omega$ is closed. In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) for all $x^{*}$. In our analysis where this additional assumption does not hold, Theorem 2 only ensures that $\gamma(x^{*})$ is a generalized local solution, as the subsequent claim ensuring that $\gamma(x^{*})$ is a (usual) local solution relies on a sufficient but not necessary condition. Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed.